find an equation of the line contaning the given pair of points
(-7,-5) and (-2,-1) the equation of the line in slop intercept from is y=?

Answer by G Nana(3)

m= Y2-Y1/X2-X1
-1-(-5)/-2-(-7)= 4/5 Which is the slope.
by the formula
Y-Y1=m(X-X1)
Y-(-5)= 4/5(X-(-7))
Y+5=4/5x+28/5
Y=4/5x+28/5- 5
Y=4/5x+28/5-25/5
Y=4/5x+3/5